OpenStudy (anonymous):

max worked 3 more then twice as many hours gary did. How many hours did each work if together they worked for 48 hours?

ILovePuppiesLol (ilovepuppieslol):

max

rebeccaxhawaii (rebeccaxhawaii):

2x+3

OpenStudy (anonymous):

o plz hurry

ILovePuppiesLol (ilovepuppieslol):

^

OpenStudy (anonymous):

2x + 3?

OpenStudy (anonymous):

2x + 3 = 48?

ILovePuppiesLol (ilovepuppieslol):

yes

ILovePuppiesLol (ilovepuppieslol):

wait

ILovePuppiesLol (ilovepuppieslol):

no

rebeccaxhawaii (rebeccaxhawaii):

m+g=48 2g+3=m

OpenStudy (anonymous):

so

rebeccaxhawaii (rebeccaxhawaii):

max=m g=gary

OpenStudy (anonymous):

2g + 3 = m?

OpenStudy (anonymous):

?

rebeccaxhawaii (rebeccaxhawaii):

2 x gary plus 3 = max

OpenStudy (anonymous):

but whats the total ammount of hours for max?

rebeccaxhawaii (rebeccaxhawaii):